# Identifying preferred routes of sharing information on social networks

### v1

@@ 1 Introduction
The flow of information between users influences not only their knowledge but also their decision-making processes and judgment. Today, a large part of this information flow is carried out through social networks, which have become an inseparable part of daily life. From social marketing [yang2010study] and preventing the spread of misinformation [gausen2021can], to predicting political elections [hu2017competing, lang2022opinion, lewis2005election] and natural disaster alarm [kim2018social, zhang2016comprehensive], understanding the mechanisms of information propagation within cyberspace has consequently become an important area of multidisciplinary research. Although users can theoretically connect on most contemporary online social platforms, these connections are far from uniform. Pre-existing offline social networks are the primary source of this heterogeneity, largely influenced by the geographical distribution of users [grabowicz2014entangling, crandall2010inferring]. However, additional factors, such as the online visibility of certain users (e.g., celebrities, political figures, or media companies) or shared interests that connect individuals regardless of their offline proximity or relationship, can create differences between cyberspace networks and their offline counterparts. Moreover, in both networks, connections are utilized with varying frequency depending on the topic of communication. A person may be more likely to share messages or engage with certain individuals on specific subjects while turning to different contacts for other subjects. This tendency forms a set of preferences in content sharing for each user, referred to as the underlying diffusion network [poux2023dirichlet], which are considered to be topic-dependent. The news spread can then be modeled as a biased random walk on this topic-sensitive underlying network [molaei2018information].
Understanding user preferences can help predict how news spreads on a given social networking platform, enabling applications such as targeted marketing and fake news prevention. Numerous studies have investigated various aspects of this phenomena [kurka2015online], ranging from theoretical modeling [li2017survey] and measuring [PhysRevE.78.065102] information diffusion, to implementing computational and machine learning methods that could be used for practical use [firdaus2021retweet, bunyamin2016comparison]. However, the literature reveals a notable gap in studies focused on modeling the emergence of user preferences. In this study, following a review of the phenomenology of news sharing, we propose that the formation of individual preferences can be governed by two distinct dynamics, which we refer to as the global and local preferential models. These models are inspired by the seminal work of Barabási and Albert [barabasi1999emergence], which will generate an underlying network that contains sharing preferences. Sharing of information is then modeled as a biased random walk on this underlying network. We then incorporate two measures, namely a modified weighted Jaccard index and functional similarity, to evaluate the presence of user preferences in news propagation over the networks generated by these models. As a case study, we apply our framework to political hashtags from the X platform (formerly Twitter) to demonstrate and validate our approach.
@@
Figure 1: One case of preferred routes in nature is riverbeds[pixabay_swamp], which form over time after many repetitions. The more water streams down a path, the more indented the riverbed will become. We take this idea to explain the formation of preferred routes of information sharing on social media.
@@ 2 Mathematical Model
To construct the model, we begin by examining empirical studies on the motivations and dynamics of social interactions. Rather than capturing the full behavioral complexity, we identify and incorporate key mechanisms that are most relevant for formal modeling. Social and group living has long been recognized as a survival strategy among many species, including humans, providing advantages in reproduction, offspring care, foraging, and predator avoidance [rubenstein1978predation]. In the modern context, however, the motivations for human social interaction have evolved beyond basic survival and reflect higher-order psychological needs. This shift is particularly evident in the use of social media. Wong et al. [wong2017motivations] argue that content sharing on social networks is primarily driven by the desire to inform and entertain an audience but also stems from deeper motivations such as maintaining relationships, influencing others, asserting individuality, and affiliating with social groups. According to their findings, social media engagement is largely shaped by an intrinsic need to share. They also mention that sharing content is influenced by three additional factors: (i) the recipient, (ii) the relationship between them, and (iii) the content, all of which we aim to incorporate into our model. The recipient and the strength of the relationship between the recipient and the sender determine the frequency—or equivalently, the probability—by which the content is shared. The strength of this relationship can arise from multiple causes, developing over time either internally through social media interactions or externally through offline connections.
@@
We aim to model the former mechanism using our proposed framework, while treating the latter as a predefined input parameter, since external social connections often form long before an individual is eligible to use social media. Although these external relationships can evolve after joining a platform, we assume such changes are relatively slow and can be neglected in our dynamic formulation. Overall, the key point is that these relationships are both user-specific and time-dependent. As highlighted in the third point above, the strength of a relationship can vary depending on the content being shared. According to Berger et al. [berger2012makes], sharing practically useful information is motivated either by altruistic intentions, such as helping others, or by the desire to enhance one’s social image, for example, to appear knowledgeable to colleagues. Picone et al. [picone2020shares] reinterpret this latter motivation as a form of self-publication, drawing an analogy to journalism in which individuals prioritize the interests of their audience and selectively share topics aligned with those interests. This perspective supports the assumption that each social media user has context-dependent sharing preferences. For instance, a user may exhibit distinct sharing behaviors depending on whether the content is political, entertainment-related, or scientific. Consequently, the network of shared content can be decomposed into topical multiplexes that evolve independently and simultaneously. This concept is known as thematic multiplex [hanteer2019innovative].
@@
Lastly, we also take into consideration the fact that some users are more prominent due to various reasons, such as being a celebrity or a political figure. Such prominence increases their visibility within the network and makes them more likely to be involved in social interactions [wasserman1994social]. It is important to clarify that when referring to the roles of sender and recipient, we mean the temporal sequence in which individuals are exposed to a given piece of content. These roles can manifest differently across platforms. For instance, in messaging systems like email or direct messaging features of platforms such as Instagram, the sender possesses the content before the recipient and decides whether or not to share it. In contrast, on microblogging platforms such as X, the user who originally posts the content (e.g., the one being retweeted) is the sender, and the one who retweets it is the recipient, even though in this case, it is the recipient who actively decides to share it.
@@ 2.1 Global Preference Evolution Dynamics
We consider a base network $G$, modeled as a weighted, undirected graph of size $N$, where nodes represent users. The degree $k$ of a node quantifies its prominence, and the edge weight $L$ between two nodes, an integer, represents the number of social interactions (termed events) between them. Following the framework of Rabbani et al. [rabbani2021memory], the dynamics proceed by selecting a sender node at random to initiate an event, capturing user activity. The sender is more likely to interact with a recipient of higher node degree, reflecting the increased prominence and hence visibility of such users. Once the recipient is chosen, the interaction is recorded by incrementing the corresponding edge weight by $x$, which is typically set to one when following the definition of the number of interactions, but it can also be set to other values to emphasize the influence of interactions. Accordingly, the probability of a node participating in an event, either as the sender or the recipient, at time $t_{n}$ is:
@@
This result indicates that the rate of increase in a link’s weight is directly proportional to the average prominence of its two end nodes.
@@ 2.2 Local Preference Evolution Dynamics
Due to the nature of this dynamic, which dictates that the evolution of each link’s weight is proportional to itself, all links with zero weight will remain zero indefinitely. This is inconsistent with real-world social networks, where any two individuals can potentially connect, albeit mostly with low probability. To address this, we add a small constant $L_{0}$ to all possible links in the network, ensuring weak but nonzero interaction probabilities. This approach resembles the modification introduced by Price et al. [price1965networks], a precursor to the BA model. We choose $L_{0}\ll x$, where $x$ the link weight increment is (e.g., for $x=1$, we set $L_{0}=0.01$). Consequently, the network becomes effectively fully connected. For notational simplicity, we redefine the adjusted link weights as $L_{ij}^{{}^{\prime}}:=L_{ij}+L_{0}$ for all $i,j\in G$. This leads to the following rate of change for the link weight in the continuum limit:
@@
(a) The probability of traversing the undirected link A-B when the sender is node A is approximately 37%. Here, the preferred choice of node A is node B and not node C.
(b) The probability of traversing the undirected link A-B when the sender is node B is approximately 19%. Here, the preferred choice of node B is node C and not node A.
Figure 2: The weight of the undirected link A-B can render different preferences depending on the sender node, hence we are able to model a phenomenon that is inherently directed using an undirected network.
@@ 3 Route Preference In Information Flow
We propose that the evolution of social networks can follow a mechanism incorporating both the global and local dynamics introduced in Sections 2.1 and 2.2. However, it is not possible to directly verify this through comparing the networks that the models generate with the real-world underlying preference network for direct validation of this hypothesis, as the latter is not accessible to us; at most, we can acquire a part of it (e.g. the follower-followee network, or friendship network). But even this will not suffice; indeed, for most opinion formation social interactions, social threads often involve people who don’t follow each other [kwak2010twitter]. Furthermore, studies have shown that additional factors, including the writing style and sentiment of shared content, significantly influence interaction patterns [jimenez2021sentiments, stieglitz2013emotions]. As a result, the follower-followee structure becomes a poor predictor of the likelihood of interaction. For example, on the X platform (formerly Twitter), only approximately 10% of retweet interactions involving hashtags occur between users who are connected through a follower-follower relationship [bastos2012sticks]. Instead, what is readily observable is the flow of events between users, which we will acquire from the real world social networks.
@@
3.1 Real World Data Retrieval
As an example case, we used the hashtag retweet data from the Farsi X with political contexts[mohammadikian2025retweet]. These data were acquired by Mohammadi et al. [mohammadi2022footprint] and further studied by Bigdeli et al. [bigdeli2024bots], spanning from April 29, 2021, to June 24, 2021 —7 weeks prior to Iran’s 2021 presidential election. The trending hashtags were tracked daily, and 16 of them were used in this study. For inclusiveness, the hashtags chosen consisted of those specific to each of the two opposing parties and also included neutral hashtags used by both groups. Data retrieval was done such that every tweet containing a specific hashtag from a specified date was stored in an SQLite database, accessed through the X’s Standard Search API. More specifically, each tweet and retweet was saved along with the user ID of the retweeter and the user ID of the user being retweeted, along with the hashtags used in the body of the retweet and a timestamp of the retweet happening. The number of distinct retweets containing at least one of the election hashtags was 5701902, and they were posted by 140638 unique users collected through X platform’s data, which at the time of collection was named Twitter.
A weighted, undirected retweet network was constructed from the data gathered for each hashtag, where nodes are the users and the link weights are the number of times a tweet containing that Hashtag has been shared between two users, in either directions. This retweet network can be seen as the trajectory of a biased random walk on the underlying network of preferences discussed previously in this section. Note that the choice of an undirected retweet network will not pose any problem for our analysis, as discussed in section 2.2. To further study the existence of preference, the two used measures will require us to perform pre-processing on the raw retweet data, the details of which are presented in Section 3.2.3.
@@ 3.2.1 Modified weighted Jaccard Index
Similar to the desire paths already introduced in the social sciences [nichols2014social], we would like to study the overlap of the imprint of social interactions between social network users. If there exists preferred routes for the dissemination of news of a certain topic, then all hashtags relating to that topic will follow that route more or less. If the tweet is considered as a biased random walker, the footprints of the walkers corresponding to each hashtag must have a considerable overlap with those of another hashtag. Due to the inherent probabilistic nature of news dissemination, some randomness is inevitable, and non-preferred links will also be traversed. However, what is critical is the relative frequency with which preferred links are traversed as opposed to non-preferred ones. To identify and compare the preferred dissemination routes for a given topic, we quantify how often each link associated with a hashtag is traversed (i.e., the number of times news is shared through that link), as illustrated in Fig. 3.
The typical [unweighted] Jaccard index gives the ratio of the overlap of two sets to the union of them (i.e. $J(A,B)=|A\cap B|/|A\cup B|$), which has already been a staple in studying social networks[evkoski2021community, jafariasbagh2014clustering]. Naturally, no repetition is considered if we want to calculate the Jaccard index of two multisets (a.k.a. sets but with arbitrary repetition of elements). Therefore, the unweighted Jaccard index will only take into account the links traversed without considering the number of repetitions on that link. The weighted Jaccard index [costa2021further] is a solution to measuring the overlap of multisets, where the frequency of the news sharing between each pair of users is taken into account. In order to embed the intuition behind the desire paths and the depth of footprints that retweets leave behind, we use a modification of this measure, which essentially gauges the percentage of all the steps taken in the sharing of two hashtags that have traversed mutual links. For two hashtags with adjacency matrices $\mathbf{M_{1}}$ and $\mathbf{M_{2}}$, this could be calculated as follows:

### v2

@@ 1 Introduction
The flow of information between users influences not only their knowledge but also their decision-making processes and judgment. Today, a substantial portion of this information flow is carried out through online social networks, which have become an inseparable part of daily life. From social marketing [53] and misinformation mitigation [15], to predicting political elections [22, 29, 30] and issuing natural disaster alerts [26, 55], understanding the mechanisms of information propagation in cyberspace has emerged as a central problem in multidisciplinary research.
Although users can theoretically connect with many others on contemporary online platforms, these connections are far from uniform. Pre-existing offline social structures, largely influenced by geographic constraints [19, 11], online visibility (e.g., celebrities, political figures or media organizations), and shared interests all contribute to heterogeneous interaction patterns. Moreover, connections are not used uniformly across topics. A person may be more likely to share content with certain contacts on specific subjects while engaging different contacts for other topics. This behavior gives rise to topic-dependent sharing preferences for each user, forming what is often referred to as an underlying diffusion network [43]. News propagation can then be conceptualized as a biased random walk on this topic-sensitive network [36]. In this work, we investigate whether such route-level sharing preferences can emerge from simple generative mechanisms and whether these mechanisms can reproduce empirical patterns observed on social media platforms.
A large body of research has examined information diffusion in social networks [27], including theoretical modeling approaches [31], empirical measurements of diffusion dynamics [18], and computational or machine learning methods aimed at prediction and practical applications [14, 8]. Diffusion models take the underlying network as given and model a dynamical process unfolding on top of it, such as Threshold models [20, 25, 9], Cascade models [17, 25] and many others [48], each with a property best fit to simulate a particular spreading scenario in social networks. These frameworks focus on transmission events, temporal excitation, and cascade statistics, while typically assuming fixed edge weights or homogeneous transmission rules.
While diffusion models address the process of propagation, evolution models address the structure that constrains such propagation. For example, foundational models such as the Barabási–Albert (BA) model [3] explain the emergence of hubs through a node-level reinforcement mechanism, i.e. the preferential attachment. Extensions such as the weighted growth model of Barrat, Barthélemy, and Vespignani (BBV) [4] further incorporate edge weights, allowing link strengths to evolve as the network expands. These models belong to a class of structural evolution frameworks in which the network’s growth is the primary object of study. Bridging the gap between purely dynamical models and purely evolutionary models are the co-evolving models of networks, such as the COEVOLVE [13] in which a diffusion process and link creation happens simultaneously.
Although these evolution models successfully capture node-level prominence, the literature reveals a notable gap in studies and modeling the emergence of link-level preferences and the formation and strengthening of specific ties through repeated interaction remains comparatively underexplored, which is one way to study route-level sharing preferences in social networks.
To address this gap, we propose two generative models of link-level preference formation: a global preferential model and a local preferential model. Both are inspired by the reinforcement principle underlying the BA model but extend it from node-level to link-level dynamics. While the BA mechanism produces recurrently attached nodes, our models also generate recurrently traversed links, leading to the emergence of preferred sharing paths. In both models, probabilistic reinforcement at each time step gradually shapes an underlying weighted network that encodes user preferences. Information diffusion is then modeled as a biased random walk on this emergent network. To evaluate whether the models reproduce empirical sharing patterns, we employ two complementary preference measures: a modified weighted Jaccard index and a functional similarity metric. The aforementioned BBV model and a random (no-preference) model are taken as our benchmark and null models respectively. Applying our framework to political hashtag data from the X platform (formerly Twitter), we demonstrate that link-level reinforcement-based mechanisms can generate structured sharing patterns that are substantially closer to the empirical data, compared to the benchmark and the null models.
The remainder of the paper is organized as follows. Section 2 introduces the global and local preferential models. In Section 3, we describe the dataset and preprocessing steps, followed by introducing preference measures on networks. Section 4 reports the results of preference measures on both the empirical and simulation data, followed by conclusion.
@@
Figure 1: One case of preferred routes in nature is riverbeds [41], which form over time after many repetitions. The more water streams down a path, the more indented the riverbed will become. We take this idea to explain the formation of preferred routes of information sharing on social media.
@@ 2 Mathematical Model
To construct the model, we begin by examining empirical studies on the motivations and dynamics of social interactions. Rather than capturing the full behavioral complexity, we identify and incorporate key mechanisms that are most relevant for formal modeling. Social and group living have long been recognized as a survival strategy among many species, including humans, providing advantages in reproduction, offspring care, foraging, and predator avoidance [47]. In the modern context, however, the motivations for human social interaction have evolved beyond basic survival and reflect higher-order psychological needs. This shift is particularly evident in the use of social media. Wong et al. [52] argue that content sharing on social networks is primarily driven by the desire to inform and entertain an audience but also stems from deeper motivations such as maintaining relationships, influencing others, asserting individuality, and affiliating with social groups. According to their findings, social media engagement is largely shaped by an intrinsic need to share. They also mention that sharing content is influenced by three additional factors: (i) the recipient, (ii) the relationship between them, and (iii) the content, all of which we aim to incorporate into our model. The recipient and the strength of the relationship between the recipient and the sender determine the frequency—or equivalently, the probability—by which the content is shared. The strength of this relationship can arise from multiple causes, developing over time either internally through social media interactions or externally through offline connections.
@@
We aim to model the former mechanism using our proposed framework, while treating the latter as a predefined input parameter, since external social connections often form long before an individual is eligible to use social media. Although these external relationships can evolve after joining a platform, we assume such changes are relatively slow and can be neglected in our dynamic formulation. Overall, the key point is that these relationships are both user-specific and time-dependent. As highlighted in the third point above, the strength of a relationship can vary depending on the content being shared. According to Berger et al. [6], sharing practically useful information is motivated either by altruistic intentions, such as helping others, or by the desire to enhance one’s social image, for example, to appear knowledgeable to colleagues. Picone et al. [42] reinterpret this latter motivation as a form of self-publication, drawing an analogy to journalism in which individuals prioritize the interests of their audience and selectively share topics aligned with those interests. This perspective supports the assumption that each social media user has context-dependent sharing preferences. For instance, a user may exhibit distinct sharing behaviors depending on whether the content is political, entertainment-related, or scientific. Consequently, the network of shared content can be decomposed into topical multiplexes that evolve independently and simultaneously. This concept is known as thematic multiplex [21].
@@
Lastly, we also take into consideration the fact that some users are more prominent due to various reasons, such as being a celebrity or a political figure. Such prominence increases their visibility within the network and makes them more likely to be involved in social interactions [51]. It is important to clarify that when referring to the roles of sender and recipient, we mean the temporal sequence in which individuals are exposed to a given piece of content. These roles can manifest differently across platforms. For instance, in messaging systems like email or direct messaging features of platforms such as Instagram, the sender possesses the content before the recipient and decides whether or not to share it. In contrast, on microblogging platforms such as X, the user who originally posts the content (e.g., the one being retweeted) is the sender, and the one who retweets it is the recipient, even though in this case, it is the recipient who actively decides to share it.
@@ 2.1 Global Preference Evolution Dynamics
We consider a base network $G$, modeled as a weighted, undirected graph of size $N$, where nodes represent users. The degree $k$ of a node quantifies its prominence, and the edge weight $L$ between two nodes, an integer, represents the number of social interactions (termed events) between them. Following the framework of Rabbani et al. [45], the dynamics proceed by selecting a sender node at random to initiate an event, capturing user activity. The sender is more likely to interact with a recipient of higher node degree, reflecting the increased prominence and hence visibility of such users. Once the recipient is chosen, the interaction is recorded by incrementing the corresponding edge weight by $x$, which is typically set to one when following the definition of the number of interactions, but it can also be set to other values to emphasize the influence of interactions. Accordingly, the probability of a node participating in an event, either as the sender or the recipient, at time $t_{n}$ is:
@@
This result is also consistent with the fact that $k_{i}=\sum\limits_{j=1}^{N}L_{ij}$ and indicates that the rate of increase in a link’s weight is directly proportional to the average prominence of its two end nodes.
@@ 2.2 Local Preference Evolution Dynamics
Due to the nature of this dynamic, which dictates that the evolution of each link’s weight is proportional to itself, all links with zero weight will remain zero indefinitely. This is inconsistent with real-world social networks, where any two individuals can potentially connect, albeit mostly with low probability. To address this, we add a small constant $L_{0}$ to all possible links in the network, ensuring weak but nonzero interaction probabilities. This approach resembles the modification introduced by Price et al. [44], a precursor to the BA model. We choose $L_{0}\ll x$, where $x$ is the link weight increment (e.g., for $x=1$, we set $L_{0}=0.01$). Consequently, the network becomes effectively fully connected. For notational simplicity, we redefine the adjusted link weights as $L_{ij}^{{}^{\prime}}:=L_{ij}+L_{0}$ for all $i,j\in G$. This leads to the following rate of change for the link weight in the continuum limit:
@@
(a) The probability of traversing the undirected link A-B (highlighted blue) when the sender is node A is approximately 37%. Here, the preferred choice of node A is node B (colored green)
and not node C.
(b) The probability of traversing the undirected link A-B (highlighted blue) when the sender is node B is approximately 19%. Here, the preferred choice of node B is node C (colored green)
and not node A.
Figure 2: A model with undirected network can account for an inherently directed phenomenon. For instance, the weight of the undirected link A-B can render different preferences depending on the sender node, breaking the directional symmetry of the role of an undirected link in news diffusion, while offering computational efficiency.
@@ 3 Route Preference In Information Flow
We propose that the evolution of social networks can follow a mechanism incorporating both the global and local dynamics introduced in Sections 2.1 and 2.2. However, it is not possible to directly verify this through comparing the networks that the models generate with the real-world underlying preference network for direct validation of this hypothesis, as the latter is not accessible to us; at most, we can acquire a part of it (e.g. the follower-followee network, or friendship network). But even this will not suffice; indeed, for most opinion formation social interactions, social threads often involve people who don’t follow each other [28]. Furthermore, studies have shown that additional factors, including the writing style and sentiment of shared content, significantly influence interaction patterns [24, 49]. As a result, the follower-followee structure becomes a poor predictor of the likelihood of interaction. For example, on the X platform (formerly Twitter), only approximately 10% of retweet interactions involving hashtags occur between users who are connected through a follower-follower relationship [5]. Instead, what is readily observable is the flow of events between users, which we will acquire from the real-world social networks.
@@
3.1 Real-World Data Retrieval
As an example case, we used the hashtag retweet data from the Farsi X within political contexts[35]. These data were acquired by Mohammadi et al. [34] and further studied by Bigdeli et al. [7], spanning from April 29, 2021, to June 24, 2021 (7 weeks prior to Iran’s 2021 presidential election). The trending hashtags were tracked daily, and 16 of them were used in this study. For inclusiveness and in order to minimize selection bias, the hashtags chosen consisted of those specific to each of the two opposing parties and also included neutral hashtags used by both groups. To collect the data, a custom program called Twitter Machine was developed (named as such since at the time of data retrieval, X platform was officially called Twitter). Given a hashtag from the collected list and a start date, the program retrieved all tweets containing the specified hashtag via the Twitter Standard Search API and stored them in an SQLite database. The application parsed the JSON objects returned by the API, extracted the relevant metadata, and updated a user index maintained in a PostgreSQL relational database. While continuous automated requests were implemented to minimize data loss, completeness of retrieval is inherently subject to the technical constraints and rate limits of the Twitter Standard Search API. All users participating in election-related discussions identified by Twitter Machine were recorded in PostgreSQL to enable relational analyses. To ensure uninterrupted and efficient data collection, the program was deployed on a Virtual Private Server (VPS), allowing automated and continuous API requests within the platform’s rate limits. The collected data were securely transferred from the VPS to local storage infrastructure, where they were archived and processed for subsequent analysis. Each tweet and retweet was saved along with the user ID of the retweeter and the user ID of the user being retweeted, along with the hashtags used in the body of the retweet and a timestamp of the retweet happening. The number of distinct retweets containing at least one of the election hashtags was 5701902, and they were posted by 140638 unique users.
A weighted, undirected retweet network was constructed from the data gathered for each hashtag, where nodes are the users and the link weights are the number of times a tweet containing that hashtag has been shared between two users, in either direction. This retweet network can be seen as the trajectory of a biased random walk on the underlying network of preferences discussed previously in this section. Note that the choice of an undirected retweet network will not pose any problem for our analysis, as discussed in Section 2.2. To further study the existence of preference, the two used measures will require us to perform preprocessing steps on the raw retweet data, the details of which are presented in Section 3.2.3.
@@ 3.2.1 Modified weighted Jaccard Index
Similar to the desire paths already introduced in the social sciences [40], we would like to study the overlap of the imprint of social interactions between social network users. If there exists preferred routes for the dissemination of news of a certain topic, then all hashtags relating to that topic will follow that route more or less. If the tweet is considered a biased random walker, the footprints of the walkers corresponding to each hashtag must have a considerable overlap with those of another hashtag. Due to the inherent probabilistic nature of news dissemination, some randomness is inevitable, and non-preferred links will also be traversed. However, what is critical is the relative frequency with which preferred links are traversed as opposed to non-preferred ones. To identify and compare the preferred dissemination routes for a given topic, we quantify how often each link associated with a hashtag is traversed (i.e., the number of times news is shared through that link), as illustrated in Fig. 3.
The typical [unweighted] Jaccard index gives the ratio of the overlap of two sets to the union of them (i.e. $J(A,B)=|A\cap B|/|A\cup B|$), which has already been a staple in studying social networks[12, 23]. Naturally, no repetition is considered if we want to calculate the Jaccard index of two multisets (a.k.a. sets but with arbitrary repetition of elements). Therefore, the unweighted Jaccard index will only take into account the links traversed without considering the number of repetitions on that link. The weighted Jaccard index [10] is a solution to measuring the overlap of multisets, where the frequency of the news sharing between each pair of users is taken into account. In order to embed the intuition behind the desire paths and the depth of footprints that retweets leave behind, we use a modification of this measure, which essentially gauges the percentage of all the steps taken in the sharing of two hashtags that have traversed mutual links. For two hashtags with adjacency matrices $\mathbf{M_{1}}$ and $\mathbf{M_{2}}$, this could be calculated as follows: